# Neural Bridge Processes

### v1

@@ A.1 Neural Processes and Their Extensions
Neural Processes (NPs) Garnelo et al. (2018b) combine the flexibility of neural networks with the uncertainty modeling capabilities of stochastic processes, aiming to learn distributions over functions. Conditional Neural Processes (CNPs) Garnelo et al. (2018a) extend this framework by conditioning on observed context points to predict target outputs. Attentive Neural Processes (ANPs) Kim et al. (2019) further enhance NPs by incorporating attention mechanisms, enabling the model to focus on relevant context points for each target prediction. Despite these advancements, challenges remain in capturing complex, multimodal distributions and ensuring consistency in posterior predictions. These approaches have been successfully extended to various domains, including sequential modeling Singh et al. (2019); Nguyen & Grover (2022); Bruinsma et al. (2023), convolutional architectures Gordon et al. (2019); Foong et al. (2020), graph-based models Hu et al. (2023), and probabilistic predictive models for large language models (LLMs) Requeima et al. (2024).
A.2 Diffusion and Bridge Models in Generative Modeling
Denoising Diffusion Probabilistic Models (DDPMs) Ho et al. (2020) are powerful generative models that learn data distributions by reversing a progressive noising process. Conditional Diffusion Models (CDMs) Choi et al. (2021); Zhang et al. (2023); Zhu et al. (2023) extend this framework for conditional generation by incorporating auxiliary information. Denoising Diffusion Bridge Models (DDBMs) Zhou et al. (2023); Yue et al. (2023); Zheng et al. (2024); Li et al. (2023); Peluchetti (2023); He et al. (2024); Shi et al. (2023); Naderiparizi et al. (2025) offer a natural alternative by introducing diffusion bridges—stochastic processes that interpolate between two paired distributions given as endpoints. This framework enables effective modeling of tasks such as image-to-image translation. However, existing DDBMs primarily focus on transformations in data space and may fall short in fully capturing the stochastic nature of functional mappings. In this work, we avoid the complexity and computational overhead of SDE-based diffusion bridges and instead extend the bridge concept within the DDPM Ho et al. (2020) framework through SNR-aware functional modeling.
@@ A.3 Generative Models for Function Modeling
Recent work has explored the use of diffusion models for function modeling. Neural Diffusion Processes (NDPs) Dutordoir et al. (2023) model distributions over functions by applying diffusion processes in latent space, allowing the representation of complex, non-Gaussian function distributions. Geometric Neural Diffusion Processes Mathieu et al. (2023) further extend this approach by incorporating geometric priors for infinite-dimensional modeling in non-Euclidean spaces. In parallel, other generative modeling techniques such as Neural ODEs Chen et al. (2018); Norcliffe et al. (2021), flow matching Lipman et al. (2022); Hamad & Rosenbaum , and score-based SDE methods Song et al. (2020); Dou et al. (2025) are also being integrated into the Neural Processes (NP) framework to enhance function modeling capabilities.
Appendix B Formulation of the DBP Framework
B.1 Derivation of Equation (10) in the Main Text
In our setting, the forward process also depends on $x$. Assume that $y_{0}$ is the initial state, which may correspond to $y_{0}$ or another variable. For clarity, we assume $y_{0}$ is the initial state and $x$ is the target.
@@ B.1 Derivation of Equation (10) in the Main Text
$\text{noise terms}=\sum_{s=1}^{t}\left(\sqrt{\beta_{s}}\prod_{k=s+1}^{t}\sqrt{1-\beta_{k}}\right)\epsilon_{s}.$
@@
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\beta_{s}\prod_{k=s+1}^{t}(1-\beta_{k}).$
@@
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\beta_{s}\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}.$
@@
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\left(1-\frac{\bar{\alpha}_{s}}{\bar{\alpha}_{s-1}}\right)\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}.$
@@
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s-1}}\right).$
@@
$\text{Var}(\text{noise terms})=\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{1}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{0}}\right)+\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{2}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{1}}\right)+\cdots+\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}\right).$
@@
$\text{Var}(\text{noise terms})=\bar{\alpha}_{t}\left(\frac{1}{\bar{\alpha}_{t}}-\frac{1}{\bar{\alpha}_{0}}\right).$
@@
$\text{Var}(\text{noise terms})=1-\bar{\alpha}_{t}.$
The variance of the accumulated noise can be computed similarly to the DDPM framework and is assumed to be: $(1-\bar{\alpha}_{t})I$. Thus, Equation (10) in the main text has been proven.
B.2 Derivation of Equation (16) and (17) in the Main Text
B.2.1 The reverse process posterior
The marginal distribution $q(y_{t-1}|y_{0},x)$ is:
$q(y_{t-1}|y_{0},x)=\mathcal{N}\left(\sqrt{\bar{\alpha}_{t-1}}y_{0}+\bar{\gamma}_{t-1}x,(1-\bar{\alpha}_{t-1})I\right),$ (29)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\bar{\gamma}_{t}=\sum_{s=1}^{t}\gamma_{s}\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}}$.
The reverse process posterior is:
$q(y_{t-1}|y_{t},y_{0},x)\propto q(y_{t}|y_{t-1},x)q(y_{t-1}|y_{0},x).$ (30)
This is a product of two Gaussians:
$\mathcal{N}(y_{t};Ay_{t-1}+b,\sigma_{1}^{2}I)\times\mathcal{N}(y_{t-1};\mu,\sigma_{2}^{2}I),$ (31)
where:
$A=\sqrt{1-\beta_{t}}$,
$b=\gamma_{t}x$,
$\sigma_{1}^{2}=\beta_{t}$,
$\mu=\sqrt{\bar{\alpha}_{t-1}}y_{0}+\bar{\gamma}_{t-1}x$,
$\sigma_{2}^{2}=1-\bar{\alpha}_{t-1}$.
Assume
$q(y_{t-1}|y_{t},y_{0},x)=\mathcal{N}(\tilde{\mu},\tilde{\beta}_{t}I)$
The mean of the product is:
$\tilde{\mu}=\left(\frac{A^{T}(y_{t}-b)}{\sigma_{1}^{2}}+\frac{\mu}{\sigma_{2}^{2}}\right)\left(\frac{A^{T}A}{\sigma_{1}^{2}}+\frac{1}{\sigma_{2}^{2}}\right)^{-1}.$ (32)
Substituting the values:
$\tilde{\mu}=\left(\frac{\sqrt{1-\beta_{t}}(y_{t}-\gamma_{t}x)}{\beta_{t}}+\frac{\sqrt{\bar{\alpha}_{t-1}}y_{0}+\bar{\gamma}_{t-1}x}{1-\bar{\alpha}_{t-1}}\right)\left(\frac{1-\beta_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}\right)^{-1}.$ (33)
Simplify the denominator:
$\frac{1-\beta_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}=\frac{(1-\beta_{t})(1-\bar{\alpha}_{t-1})+\beta_{t}}{\beta_{t}(1-\bar{\alpha}_{t-1})}=\frac{1-\bar{\alpha}_{t}}{\beta_{t}(1-\bar{\alpha}_{t-1})},$ (34)
since $\bar{\alpha}_{t}=(1-\beta_{t})\bar{\alpha}_{t-1}$.
Thus:
$\tilde{\mu}=\left(\frac{\sqrt{1-\beta_{t}}(y_{t}-\gamma_{t}x)}{\beta_{t}}+\frac{\sqrt{\bar{\alpha}_{t-1}}y_{0}+\bar{\gamma}_{t-1}x}{1-\bar{\alpha}_{t-1}}\right)\frac{\beta_{t}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}.$ (35)
Now, expand the numerator:
$\tilde{\mu}=\frac{\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}(y_{t}-\gamma_{t}x)+\frac{\beta_{t}(\sqrt{\bar{\alpha}_{t-1}}y_{0}+\bar{\gamma}_{t-1}x)}{1-\bar{\alpha}_{t}}.$ (36)
This can be rewritten as:
$\tilde{\mu}=\frac{\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}y_{t}+\frac{\beta_{t}\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t}}y_{0}+\left(\frac{\beta_{t}\bar{\gamma}_{t-1}}{1-\bar{\alpha}_{t}}-\frac{\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t}}{1-\bar{\alpha}_{t}}\right)x.$ (37)
The variance $\tilde{\beta}_{t}$ is derived as:
$\tilde{\beta}_{t}=\left(\frac{1-\beta_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}\right)^{-1}=\frac{\beta_{t}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}.$ (38)
B.2.2 Reparameterization of $y_{0}$ (Similar to DDPM)
From the forward process, we can express $y_{t}$ as:
$y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (39)
where $\epsilon\sim\mathcal{N}(0,I)$. Solving for $y_{0}$:
@@ B.2.2 Reparameterization of $y_{0}$ (Similar to DDPM)
Substitute $y_{0}$ into the Reverse Process Mean $\tilde{\mu}$
The derived mean $\tilde{\mu}$ is:
$\tilde{\mu}=\frac{\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})y_{t}+\beta_{t}\sqrt{\bar{\alpha}_{t-1}}y_{0}+(\beta_{t}\bar{\gamma}_{t-1}-\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t})x}{1-\bar{\alpha}_{t}}.$ (41)
Substitute $y_{0}$:
$\tilde{\mu}=\frac{\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})y_{t}+\beta_{t}\sqrt{\bar{\alpha}_{t-1}}\left(\frac{y_{t}-\bar{\gamma}_{t}x-\sqrt{1-\bar{\alpha}_{t}}\epsilon}{\sqrt{\bar{\alpha}_{t}}}\right)+(\beta_{t}\bar{\gamma}_{t-1}-\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t})x}{1-\bar{\alpha}_{t}}.$ (42)
Simplify the Expression:
1. Combine $y_{t}$ Terms:
@@
$\frac{\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})y_{t}+\frac{\beta_{t}\sqrt{\bar{\alpha}_{t-1}}y_{t}}{\sqrt{\bar{\alpha}_{t}}}}{1-\bar{\alpha}_{t}}$ $=\frac{\left(\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})+\frac{\beta_{t}}{\sqrt{1-\beta_{t}}}\right)y_{t}}{1-\bar{\alpha}_{t}},$ (43)
$=\frac{y_{t}}{\sqrt{1-\beta_{t}}}$
where we used $\bar{\alpha}_{t}=\bar{\alpha}_{t-1}(1-\beta_{t})$, so $\sqrt{\bar{\alpha}_{t-1}}/\sqrt{\bar{\alpha}_{t}}=1/\sqrt{1-\beta_{t}}$.
2. Combine $x$ Terms:
$\frac{-\frac{\beta_{t}\sqrt{\bar{\alpha}_{t-1}}\bar{\gamma}_{t}}{\sqrt{\bar{\alpha}_{t}}}+\beta_{t}\bar{\gamma}_{t-1}-\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t}}{1-\bar{\alpha}_{t}}x\overset{\triangle}{=}C(t)x$ (44)
3. Noise ($\epsilon$) Term:
$\frac{\beta_{t}\sqrt{\bar{\alpha}_{t-1}}\sqrt{1-\bar{\alpha}_{t}}\epsilon}{\sqrt{\bar{\alpha}_{t}}(1-\bar{\alpha}_{t})}=-\frac{\beta_{t}}{\sqrt{1-\beta_{t}}\sqrt{1-\bar{\alpha}_{t}}}\epsilon.$ (45)
Then, the mean can be written as:
@@ 3. Noise ($\epsilon$) Term:
$\tilde{\mu}=\frac{1}{\sqrt{1-\beta_{t}}}\left(y_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon\right)+C(t)x.$ (46)
which proves the Equation (16) in the main text.
B.2.3 Calculation of $C(t)$
From Equation (44),
$C(t)=\frac{-\frac{\beta_{t}\sqrt{\bar{\alpha}_{t-1}}\bar{\gamma}_{t}}{\sqrt{\bar{\alpha}_{t}}}+\beta_{t}\bar{\gamma}_{t-1}-\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t}}{1-\bar{\alpha}_{t}}.$ (47)
Using $\sqrt{\bar{\alpha}_{t}}=\sqrt{\bar{\alpha}_{t-1}}\sqrt{1-\beta_{t}}$, we first compute:
$\frac{\beta_{t}\sqrt{\bar{\alpha}_{t-1}}\bar{\gamma}_{t}}{\sqrt{\bar{\alpha}_{t-1}}\sqrt{1-\beta_{t}}}+\beta_{t}\bar{\gamma}_{t-1}=-\frac{\beta_{t}\bar{\gamma}_{t}}{\sqrt{1-\beta_{t}}}+\beta_{t}{\bar{\gamma}}_{t-1}.$ (48)
Next, using the relation $\bar{\gamma}_{t}=\gamma_{t}+\sqrt{1-\beta_{t}}\bar{\gamma}_{t-1}$ (since $\bar{\gamma}_{t}=\sum_{s=1}^{t}\gamma_{s}\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}}=\gamma_{t}+\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}\bar{\gamma}_{t-1}=\gamma_{t}+\sqrt{1-\beta_{t}}\bar{\gamma}_{t-1}$), we simplify:
$-\frac{\beta_{t}\bar{\gamma}_{t}}{\sqrt{1-\beta_{t}}}+\beta_{t}\bar{\gamma}_{t-1}$ $=-\frac{\beta_{t}(\gamma_{t}+\sqrt{1-\beta_{t}}\bar{\gamma}_{t-1})}{\sqrt{1-\beta_{t}}}+\beta_{t}\bar{\gamma}_{t-1}=-\frac{\beta_{t}\gamma_{t}}{\sqrt{1-\beta_{t}}}-\beta_{t}\bar{\gamma}_{t-1}+\beta_{t}\bar{\gamma}_{t-1}$ (49)
$=-\frac{\beta_{t}\gamma_{t}}{\sqrt{1-\beta_{t}}}.$
Thus, the final bridge correction term is:
$C_{t}(x)=\frac{-\frac{\beta_{t}\gamma_{t}}{\sqrt{1-\beta_{t}}}-\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t}}{1-\bar{\alpha}_{t}}x=-\frac{\beta_{t}+(1-\beta_{t})(1-\bar{\alpha}_{t-1})}{\sqrt{1-\beta_{t}}(1-\bar{\alpha}_{t})}\gamma_{t}x.$ (50)
Using $\sqrt{\bar{\alpha}_{t}}=\sqrt{\bar{\alpha}_{t-1}}\sqrt{1-\beta_{t}}$, we obtain the simplified form:
$C(t)x=-\frac{\gamma_{t}}{\sqrt{1-\beta_{t}}}x.$ (51)
Substitute $\gamma_{t}=\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}}=\frac{\bar{\alpha}_{T}(1-\bar{\alpha}_{t})}{\bar{\alpha}_{t}(1-\bar{\alpha}_{T})}$:
$C(t)=-\frac{1}{\sqrt{1-\beta_{t}}}\cdot\frac{\bar{\alpha}_{T}(1-\bar{\alpha}_{t})}{\bar{\alpha}_{t}(1-\bar{\alpha}_{T})}.$ (52)
In essence, the role of $C(t)$ is to correct the contribution of $x$ during the reverse process, ensuring that the generative procedure properly incorporates the bridging information.
B.3 Derivation of Equation (18) in the Main Text
@@
We derive the NBP loss $\mathcal{L}_{\theta}=\mathbb{E}_{t,y_{0},x,\epsilon}\left[\|\epsilon_{\theta}(y_{t},x,t)-\epsilon\|_{2}^{2}\right]$ in Equation (18) in the main text directly from the Evidence Lower Bound (ELBO).

### v2

@@ A.1 Neural Processes and Their Extensions
Neural Processes (NPs) [14] combine the flexibility of neural networks with the uncertainty modeling capabilities of stochastic processes, aiming to learn distributions over functions. Conditional Neural Processes (CNPs) [13] extend this framework by conditioning on observed context points to predict target outputs. Attentive Neural Processes (ANPs) [23] further enhance NPs by incorporating attention mechanisms, enabling the model to focus on relevant context points for each target prediction. Despite these advancements, challenges remain in capturing complex, multimodal distributions and ensuring consistency in posterior predictions. These approaches have been successfully extended to various domains, including sequential modeling [41, 32, 3], convolutional architectures [16, 11], graph-based models [20], and probabilistic predictive models for large language models (LLMs) [38].
A.2 Diffusion, Conditional Diffusion, and Bridge Models
Denoising Diffusion Probabilistic Models (DDPMs) [19] are powerful generative models that approximate complex data distributions by reversing a progressive noising process. Conditional Diffusion Models (CDMs) [7, 50, 54] extend this framework by incorporating auxiliary information, enabling conditional generation. In most CDMs, however, the forward noising process remains unconditional and the condition is injected into the reverse denoising network. This makes CDMs broadly applicable, but it does not explicitly define a condition-dependent forward trajectory or a condition-aligned terminal distribution. NBP differs from this design choice: the input $x$ appears directly in the forward transition kernel, so conditioning affects the entire diffusion path rather than only the learned denoiser.
More recently, Denoising Diffusion Bridge Models (DDBMs) [53, 49, 52, 25, 36, 18, 40, 31] have been proposed as a natural alternative. DDBMs introduce diffusion bridges—stochastic processes that interpolate between two paired data distributions given as endpoints—making them well-suited for tasks such as image-to-image translation. NBP shares the general bridge intuition of endpoint-aware generative paths, but its target is different. Rather than learning a generic data-space transport between two domains, NBP is designed for stochastic function modeling from context-target observations. The endpoint anchor is the function input, or its learned projection when dimensions differ, and the bridge is used to strengthen conditional function prediction. To the best of our knowledge, this is the first use of bridge-anchored generative paths as a Neural Process mechanism.
Thus, NBP should be viewed neither as a standard CDM nor as a direct application of existing DDBMs. It is a DDPM-style bridge formulation specialized to Neural Process-style function learning: it preserves the architecture and training protocol of the conditional diffusion baseline NDP while modifying the process-level forward kernel and corresponding reverse correction. Since DDBMs address a different problem setting, a direct experimental comparison with the original DDBM is not the main focus; instead, we compare against diffusion-based function modeling baselines such as NDP, Geometric NDP, and SNP, and use component ablations to rule out simpler explanations such as endpoint initialization, coordinate-anchor capacity, or a stronger denoiser.
@@ A.3 Generative Models for Function Modeling
Recent work has explored the use of diffusion models for function modeling. Neural Diffusion Processes (NDPs) [10] model distributions over functions by applying diffusion processes in latent space, allowing the representation of complex, non-Gaussian function distributions. Geometric Neural Diffusion Processes [30] further extend this approach by incorporating geometric priors for infinite-dimensional modeling in non-Euclidean spaces. In parallel, other generative modeling techniques such as Neural ODEs [6, 34], flow matching [26, 17], and score-based SDE methods [44, 8] are also being integrated into the Neural Processes (NP) framework to enhance function modeling capabilities. In our experiments, we compared against open-source methods, including score-based neural processes (SNP) [8] and Geometric Neural Diffusion Processes [30], both of which demonstrated the empirical advantages of our approach.
A.4 Flow Matching Neural Processes
Flow Matching Neural Processes (FlowNP) [17] provide a recent flow-based alternative to diffusion-style Neural Processes. Rather than learning a denoising score or noise predictor over a discrete diffusion chain, FlowNP defines a continuous conditional probability path between a simple base sample $y_{0}$ and the observed target value $y_{1}$, commonly using a linear interpolation of the form $y_{t}=(1-t)y_{0}+ty_{1}$. A transformer-based Neural Process backbone then predicts the corresponding velocity field conditioned on the context set. At test time, conditional samples are obtained by integrating the learned vector field, and likelihoods can be estimated through the ODE change-of-variables formula.
This perspective is useful for two reasons. First, FlowNP is a strong and recent baseline for stochastic function modeling, so comparisons against it help clarify whether improvements come from the bridge construction rather than from using a stronger backbone. Second, its path-based formulation is closely related to our motivation: both approaches view conditional function prediction as learning a generative trajectory over function values. The key distinction is that FlowNP uses an input-independent interpolation path and injects context through the velocity network, whereas NBP modifies the generative path itself by anchoring intermediate states to the input. In additional controlled experiments, we therefore also consider bridge-augmented flow paths using the same FlowNP architecture and likelihood evaluation protocol.
Appendix B Proofs for Neural Bridge Processes
B.1 Proof of Proposition 4.1
We derive the closed-form marginal distribution in Proposition 4.1. The forward process depends on the input anchor $x$.
@@ Appendix B Proofs for Neural Bridge Processes
$\text{noise terms}=\sum_{s=1}^{t}\left(\sqrt{\beta_{s}}\prod_{k=s+1}^{t}\sqrt{1-\beta_{k}}\right)\epsilon_{s}.$ (28)
@@
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\beta_{s}\prod_{k=s+1}^{t}(1-\beta_{k}).$ (29)
@@
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\beta_{s}\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}.$ (30)
@@
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\left(1-\frac{\bar{\alpha}_{s}}{\bar{\alpha}_{s-1}}\right)\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}.$ (31)
@@
$\text{Var}(\text{noise terms})=\sum_{s=1}^{t}\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{s-1}}\right).$ (32)
@@
$\text{Var}(\text{noise terms})=\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{1}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{0}}\right)+\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{2}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{1}}\right)+\cdots+\left(\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t}}-\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}\right).$ (33)
@@
$\text{Var}(\text{noise terms})=\bar{\alpha}_{t}\left(\frac{1}{\bar{\alpha}_{t}}-\frac{1}{\bar{\alpha}_{0}}\right).$ (34)
@@
$\text{Var}(\text{noise terms})=1-\bar{\alpha}_{t}.$ (35)
The variance of the accumulated noise is therefore $(1-\bar{\alpha}_{t})I$. Thus, Eq. (8) follows.
B.2 Proof of Proposition 4.2
We write $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. The bridge forward transition is
$q(y_{t}\mid y_{t-1},x)=\mathcal{N}\!\left(y_{t};\sqrt{\alpha_{t}}y_{t-1}+\gamma_{t}x,\beta_{t}I\right),$ (36)
and Proposition 4.1 gives
$q(y_{t-1}\mid y_{0},x)=\mathcal{N}\!\left(\sqrt{\bar{\alpha}_{t-1}}y_{0}+\bar{\gamma}_{t-1}x,(1-\bar{\alpha}_{t-1})I\right).$ (37)
The accumulated bridge coefficient satisfies the one-step recursion
$\bar{\gamma}_{t}=\gamma_{t}+\sqrt{\alpha_{t}}\bar{\gamma}_{t-1},$ (38)
which follows directly from $\bar{\gamma}_{t}=\sum_{s=1}^{t}\gamma_{s}\sqrt{\bar{\alpha}_{t}/\bar{\alpha}_{s}}$.
Gaussian posterior.
The exact bridge posterior is obtained from the product
$q(y_{t-1}\mid y_{t},y_{0},x)\propto q(y_{t}\mid y_{t-1},x)\,q(y_{t-1}\mid y_{0},x).$ (39)
Since both terms are Gaussian with isotropic covariance, the posterior is $\mathcal{N}(\tilde{\mu}_{t},\tilde{\beta}_{t}I)$, where
$\tilde{\beta}_{t}=\left(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}\right)^{-1}=\frac{\beta_{t}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}},$ (40)
and
$\tilde{\mu}_{t}=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})y_{t}\beta_{t}\sqrt{\bar{\alpha}_{t-1}}y_{0}\left(\beta_{t}\bar{\gamma}_{t-1}-\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t}\right)x}{1-\bar{\alpha}_{t}}.$ (41)
This is the standard Gaussian conditioning formula applied to the shifted linear observation $y_{t}=\sqrt{\alpha_{t}}y_{t-1}+\gamma_{t}x+\sqrt{\beta_{t}}\epsilon_{t}$.
Noise-prediction form.
Using the closed-form forward marginal,
$y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\qquad\epsilon\sim\mathcal{N}(0,I),$ (42)
we rewrite
@@
Substituting Eq. (43) into Eq. (41), the coefficients of $y_{t}$, $\epsilon$, and $x$ can be collected separately. For the $y_{t}$ coefficient,
$\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})+\beta_{t}\sqrt{\bar{\alpha}_{t-1}/\bar{\alpha}_{t}}}{1-\bar{\alpha}_{t}}$ (44)
$\quad=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})+\beta_{t}/\sqrt{\alpha_{t}}}{1-\bar{\alpha}_{t}}=\frac{1}{\sqrt{\alpha_{t}}},$
where the last equality uses $\bar{\alpha}_{t}=\alpha_{t}\bar{\alpha}_{t-1}$. The noise coefficient becomes
$\frac{\beta_{t}\sqrt{\bar{\alpha}_{t-1}}\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}(1-\bar{\alpha}_{t})}=-\frac{\beta_{t}}{\sqrt{\alpha_{t}}\sqrt{1-\bar{\alpha}_{t}}}.$ (45)
Thus the only remaining non-standard term is the coefficient of $x$:
$C_{t}=\frac{-\beta_{t}\sqrt{\bar{\alpha}_{t-1}/\bar{\alpha}_{t}}\,\bar{\gamma}_{t}+\beta_{t}\bar{\gamma}_{t-1}-\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t}}{1-\bar{\alpha}_{t}}.$ (46)
Using $\sqrt{\bar{\alpha}_{t-1}/\bar{\alpha}_{t}}=1/\sqrt{\alpha_{t}}$ and the recursion in Eq. (38),
$-\frac{\beta_{t}}{\sqrt{\alpha_{t}}}\bar{\gamma}_{t}+\beta_{t}\bar{\gamma}_{t-1}$ (47)
@@
$\quad=-\frac{\beta_{t}}{\sqrt{\alpha_{t}}}\left(\gamma_{t}+\sqrt{\alpha_{t}}\bar{\gamma}_{t-1}\right)+\beta_{t}\bar{\gamma}_{t-1}=-\frac{\beta_{t}}{\sqrt{\alpha_{t}}}\gamma_{t}.$
Therefore,
$C_{t}$ $=\frac{-\frac{\beta_{t}}{\sqrt{\alpha_{t}}}\gamma_{t}-\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\gamma_{t}}{1-\bar{\alpha}_{t}}$ (48)
$=-\frac{\gamma_{t}}{\sqrt{\alpha_{t}}}\frac{\beta_{t}+\alpha_{t}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}=-\frac{\gamma_{t}}{\sqrt{\alpha_{t}}},$
because $\beta_{t}+\alpha_{t}(1-\bar{\alpha}_{t-1})=1-\alpha_{t}\bar{\alpha}_{t-1}=1-\bar{\alpha}_{t}$.
Combining the three terms gives
@@
$\tilde{\mu}_{t}=\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon\right)-\frac{\gamma_{t}}{\sqrt{\alpha_{t}}}x.$ (49)
Replacing the true noise $\epsilon$ by the learned predictor $\epsilon_{\theta}(y_{t},x,t)$ yields the parameterized reverse mean in the main text, with
$C_{t}(x)=-\frac{\gamma_{t}}{\sqrt{\alpha_{t}}}x=-\frac{\gamma_{t}}{\sqrt{1-\beta_{t}}}x.$ (50)
This simplification does not ignore the accumulated term $\bar{\gamma}_{t}x$; it uses the exact recursion in Eq. (38), so the accumulated bridge contribution cancels with the prior mean contribution in the Gaussian posterior. The result holds for any scalar bridge schedule $\{\gamma_{t}\}_{t=1}^{T}$ used in the forward transition; the SNR schedule is one particular choice.
B.3 Proof of Proposition 4.3
@@
We derive the NBP loss in Proposition 4.3 directly from the Evidence Lower Bound (ELBO).